OpenStudy (anonymous):

Which equation models the problem below if c represents the number of puppies whose eyes are closed? There are 15 puppies in a photograph, and 11 of them have their eyes open. How many puppies have their eyes closed?

OpenStudy (anonymous):

15-11=c

OpenStudy (anonymous):

that isnt an asnwer on the list .. ( c – 11 = 15 is it this ?) 11 – c = 15 15 + c = 11 11 + c = 15